### v1

@@ 7.5. Costs of Deanonymizations
Finally, we measure the processing overhead and data transfers during the deanonymization process. Here, we look at an increasing threshold of consortium members participating in the deanonymization. For each measurement, we consider a new Seldom-anonymized circuit and start with the deanonymization request by a LEA. We repeat each measurement 42 times. We ignore network latency (negligible) and assume instantaneous decision making by humans in the loop (likely taking hours to days in real life).
We plot the time required to decrypt the layer-encrypted identities in Figure 9 (stacked bars on the left of each group). Here, the time for the first decryption layer also includes the time for the flow record. Potentially surprising, the last step in the decryption is notably faster than the previous steps. This speedup results from decryption shares being collected only by the LEA and not all consortium members. Overall, we observe a linear growth in processing with an increase in consortium members. Still, even for 75 members, processing on a single CPU core per participant requires only (95%-CI:[; ]). Thus, processing is orders of magnitude faster than human processes and does not constitute a processing bottleneck.
The other relevant performance metric is the amount of exchanged data. We thus report the total (dark blue bar) and per-consortium member (yellow bar) amount of application layer data sent and received by each party during deanonymization in Figure 9. For the exchanged data per consortium member, we observe linear growth with the number of participants. As all data is broadcasted to all participants, we observe quadratic growth in the total amount of exchanged data. Still, the total amount of exchanged data is well below and thus insignificant. Overall, we conclude that the deanonymization process is efficient and manual processes outside Seldom, such as the identification of suspicious traffic and the assessment of requests, are significantly slower.
@@ 8. Security Discussion
After showing Seldom’s remarkable performance, we discuss potential attacks against Seldom.
8.1. Oblivious Authentication Protocol
@@
Seldom enables the exit relay to learn the encrypted client identity (cf. Figure 2) and indisputably link connections to it, while preserving client anonymity (cf. Section 5). To show that this claim holds, we analyze which relays have access to which data and derive whether any relay has unindented access to client-identifying data. Here, we consider information that can be correlated within one and across different circuits.
@@
During circuit establishment, the entry relay learns the client’s IP certificate, the hop keys and , as well as the signature with . The certificate (and consequently the reuse of even across multiples circuit establishments) contains no new information about the client, as its IP address and the current time are known by the entry relay.
@@
The hop key and, consequently, the signature with are unique per circuit and derived from a secure source of randomness. Thus, they do not allow any relay to gain information about clients from other circuits. The -th middle relays knows the hop key and , the identity of the previous relay, and signatures by the previous relay and . All other information is threshold-encrypted to the relay. In Tor, intermediary relays also know their direct predecessors. As is only shared with the preceding relay and only with the following relay, non-connected relays cannot notice their inclusion in the same circuit.
@@
8.2. Collusion Within the Consortium
@@
To ensure that the risk of being covertly deanonymized in Seldom through collusion of a majority of consortium members and an exit relay is comparable to those in Tor, we assemble consortia accordingly. Here, it is important to recall that Tor does not fully protect the anonymity of its users, but only decreases the chances that an attacker can learn a client’s identity (Arma, 2014). One attack to deanonymize a client in Tor is through the collusion of the three relays composing a circuit. Thus, Seldom’s deanonymizations should be more resilient to collusions than collusion of three randomly selected relays. As the exit relay is involved in the deanonymization, we can remove it from the equation.
@@
Hence, we compare the risk of collusion in Seldom’s consortium to the risk of collusion of two randomly selected relays. There are multiple reasons why Seldom’s consortium is significantly less likely to collude than these two relays: While relays can be deployed by anyone en masse, the identities of the consortium members are publicly known. Specifically, it would be necessary to convince a significant fraction of privacy advocates, e.g., some of the directory authorities, and foreign governments to partake in such an attack. Even attempting this could result in significant repercussions, which makes a successful collusion attack among the consortium members highly unlikely. We also designed Seldom in such a way that a higher number of members must participate in a decryption than have to accept a deanonymization vote, i.e., some disagreeing members have to accept the consensus and cooperate, to further minimize the risk of collusion.
8.3. Mismatching Encrypted Identities
A malicious consortium member can use partial deanonymizations to leak information about users. Therefore, the malicious member could set up a circuit that includes the threshold-encrypted identity of a targeted user. The consortium could then be tricked to decrypt this identity, e.g., by marking the circuit as probe traffic.
However, such an attack would be detected after the consortium removes one layer of threshold encryption due to an invalid signature. The corresponding relay would be flagged as suspicious, but the malicious entity, as part of the consortium, would learn the identity of the preceding relay in the targeted circuit. The attacker could repeat this step and eventually learn the client’s identity. However, the consortium would notice the second iteration of this attack as the plaintexts are from different layers of the same circuit and thus clearly identify an attack from within the consortium. Thus, the clients identity remains secure, as long as the attacker not operated the entry relay of the targeted user. Additionally, the public would learn about this one-off attack and could abandon the network or hold the consortium member accountable.
8.4. Traffic Analysis Attacks
Traffic analysis is a common attack against anonymity networks (Murdoch and Danezis, 2005; Panchenko et al., 2011). While Tor does not explicitly protect against global adversaries (Dingledine et al., 2004), i.e., an entity that intercepts all traffic, analyzing traffic between two or multiple nodes is often possible for adversaries. For our discussion, we focus particularly on the additional information that may be leaked about the client in Seldom compared to Tor. An observer incepting the outgoing traffic of a client not using a bridge relay learns two things, the hop key of the circuit and the client’s intention to use Seldom. The hop key does not help in deanonymizing a circuit and the intention to use the network is also leaked in Tor.
As the relays in Seldom communicate through TLS channels, even during circuit establishment, no additional information is leaked besides the amount of transmitted data. However, as usually many circuits are multiplexed over a single link, this information is unlikely to give an adversary an advantage in detecting circuit extensions. In the end, such attacks can also be executed against traffic patterns of different websites (Panchenko et al., 2011, 2016; Hayes and Danezis, 2016), which is possible for both Tor and Seldom. All traffic leaving exit relays is either anonymized client traffic or encrypted deanonymization information. Neither leaks any additional information.
@@ 9. Ethical and Technical Concerns
The safety-privacy dilemma remains particularly divisive, with LEA and privacy advocates maintaining fundamentally incompatible positions regarding a problem that requires some sort of compromise (Twenning and Baier, 2024). Fostering the understanding of technical means for enabling exceptional access without the dooming vision of privacy-invasive legislation nor hidden backdoors for LEA, we address frequent concerns regarding Seldom’s design in the following.
@@
Q1 – Do we advertise for Seldom to replace Tor? No. Seldom relies on translucency and public oversight, which means that observed misbehavior by Seldom’s consortium must be met with revocations of these privileges, i.e., a migration of the user base back to Tor. In the far future, Seldom might attract a significantly larger user base than Tor and thus offer better anonymity even when accounting for selective deanonymization capabilities, especially if users start abandoning Tor. For Seldom’s security concept to work, there must, however, always exist a quick and easy way to migrate back to an anonymity network without deanonymization capabilities, even if temporarily deserted.
@@
Q2 – Can citizens trust the consortium? Our assumptions are analogous to those for secure multiparty computation: The consortium consists of entities that pursue different interests and have therefore little incentive of partaking in collusions, especially considering the high risk of such attempts being revealed to the public through the several translucency properties of Seldom. Further, Seldom protects against individual malicious consortium members. Namely, the majority of honest members will faithfully delete their secret shares after encrypted identities reach the end of the mandated retention period or in extreme events, such as an imminent government takeover. This way, the consortium renders older encrypted identities unusable for malicious members even in the case of an anticipated dishonest majority gaining control over Seldom. However, maintaining transparency about the goals and non-goals of Seldom remains important to avoid creating false expectations about Seldom’s protection.
@@
Q3 – Can individual relays misbehave? Once a relay misbehaves and does not comply fully with deanonymization requests, they will become suspicious to the requesting LEA. Hence, they would reveal themselves as targets for further investigation. To reduce the overall number of misbehaving relays, Seldom can further use active probing (cf. Section 6.4) and remove any relays that regularly misbehave.
@@
Q4 – Can investigations fail then? In exceptional cases, LEA will not be able to find the true root of harmful traffic using Seldom alone. For instance, misbehaving relays may cover up specific bad actors, or bad actors may establish circuits via compromised devices. However, all these cases lead LEA to a distinct location worth of further forensic investigation. Hence, Seldom will not always guarantee an immediate investigation success, but it provides crucial pointers to narrow down additional investigations.
@@
Q5 – Can Seldom support hidden services? Technically, Seldom can integrate hidden services. Each introduction and rendezvous point would be associated with an encrypted identity, which could be deanonymized if deemed necessary. The consortium must, however, consider the increased privacy implications of agreeing to such a request. If a Seldom-like system were to be deployed, it should however be carefully considered if hidden services should be supported or not.
@@
Q6 – Could Seldom interfere with Tor’s service? Initially, Seldom targets the population of privacy-affine users who currently shy away from using Tor due to having negative associations by tying Tor to cyber criminality. Only if Seldom is capable of establishing a reputation as the infrastructure of choice for privacy-seeking law-abiding citizens, a potential user migration from Tor to Seldom could take place. As a positive effect, this would actively harm the achievable anonymity for cybercriminals, as they cannot hide among a mass of benign traffic. On the other hand, it becomes harder to abandon Seldom if a deserted Tor networks offers little anonymity. To ensure that such a situation does not encourge extensive surveillance by the consortium, it must be easy for the user base to switch back to Tor at once, e.g., by offering a Tor and Seldom-compatible client that, by default, switches to Tor if Seldom is collectively deemed unsafe.
@@ 10. Conclusion
With this paper, we investigate the potential of selective deanonymizations for anonymity networks. We find that Seldom’s design barely impacts performance in comparision to Tor, while data overhead remains manageable. On the other hand, Seldom relies on a translucenct trusted consortium and potentially encourgages the migration of users from Tor, which would weaken the anonymity provided by Tor. With this paper we thus contribute to understanding the advantages and disadvantages of selective deanonymizations in anonymity networks like Seldom. These insights should encourage future research to mitage drawbacks, and most importantly, they should increase the objectivity in inevitable future debates about exceptional access.
@@ Appendix A Integration into Tor’s Handshake
We now discuss how Seldom integrates its oblivious authentication into Tor’s circuit establishment. Figure 10 depicts the extensions made to Tor’s circuit establishment. The data exchange of the original Tor handshake is compiled in the cell types, which are indicated in capital letters. Seldom extends the relevant cell types, i.e., Create2, Extend2, and Extended2 cells with the data necessary for the oblivious authentication. The clients and relays append the IP certificates or encrypted identities to Create2 cells. This data, however, does not always fit in a single cell. Therefore, the data is fragmented, and the first fragment is appended to the Create2 cell. All further fragments are sent in the newly added Create2_Addata cells. Seldom dispatches these cells immediately after transmitting Create2 cells. As soon as a relay receives the Create2 cell (not necessarily all Create2_Addata cells), it executes the Tor handshake and sends a Created2 cell back. After a relay sent all Create2_Addata cells, it sends a standard Extended2 cell back to the client to inform it about the successful circuit extension. Once all Create2_Addata have been received by the extending relays, they perform all checks as explained in Section 5. Overall, Seldom’s oblivious authentication handshake can thus be integrated seamlessly into Tor’s circuit extension without introducing any additional round trips.
Appendix B IP Certificates in Restricted Networks

### v2

@@ 7.5. Costs of Deanonymizations
Finally, we measure the processing overhead and data transfers during the deanonymization process. Here, we look at an increasing threshold of required consortium members participating in the deanonymization. The total size of the consortium does not influence the performance. For each measurement, we consider a new Seldom-anonymized circuit and start with the deanonymization request by an LEA. We repeat each measurement 40 times. We ignore network latency (negligible) and assume instantaneous decision-making by humans in the loop (which would likely take hours to days in real life).
We plot the time required to decrypt the layer-encrypted identities in Figure 10 (stacked bars on the left of each group). Here, the time for the first decryption layer also includes the time for the flow record. Potentially surprisingly, the last step in the decryption is notably faster than the previous steps. This speedup results from decryption shares being collected only by the LEA and not all consortium members. Overall, we observe a linear growth in processing with an increase in consortium members. Still, even for 75 members, processing on a single CPU core per participant requires only (95%-CI:[; ]). Thus, processing is orders of magnitude faster than human processes and does not constitute a processing bottleneck.
The other relevant performance metric is the amount of exchanged data. We thus report the total (dark blue bar) and per-consortium member (yellow bar) amount of application layer data sent and received by each party during deanonymization in Figure 10. For the exchanged data per consortium member, we observe linear growth with the number of participants. As all data is broadcasted to all participants, we observe quadratic growth in the total amount of exchanged data. Still, the total amount of exchanged data is well below and thus insignificant. Overall, we conclude that the deanonymization process is efficient, and manual processes outside Seldom, such as the identification of suspicious traffic and the assessment of requests, are significantly slower.
@@ 8. Security Discussion
After having evaluated its performance, we now discuss how Seldom protects against potential threats. Seldom’s threat model (cf. Section 4.2.1) discusses the two potential goals of an adversary concerning its deanonymization capabilities: (i) evade deanonymizations, possibly by misattributing suspicious traffic to an innocent client, or (ii) conduct unauthorized deanonymizations.
To analyze how Seldom thwarts attempts to bypass its deanonymization procedure, we examine the individual actors within the network. While the database operators and the consortium as a collective are considered trustworthy, relays, CAs, and individual consortium members may act maliciously (cf. Section 4.2.2). In the following, we analyze how Seldom technically prevents individuals and colluding entities from evading deanonymizations (Section 8.1) or performing unauthorized deanonymizations (Section 8.2). Finally, we discuss additional forms of collusion (Section 8.3) that lie outside the scope of Seldom’s direct technical protections.
8.1. Evading Deanonymizations
First, we look at how Seldom prevents entities from evading deanonymizations. No relay, consortium member, or client should be able to tamper with deanonymization without being clearly identified as the culprit. Malicious relays, as well as consortium members, could try to mix deanonymization data across different circuits to mount an attack. Likewise, clients could route traffic through a self-hosted relay to cover their tracks. In the following, we see how the Seldom protocol protects against such attacks.
8.1.1. Reusing Encrypted Identities
A misbehaving relay may try to hinder deanonymizations by reusing the encrypted identity of a previous client, thereby protecting the real client’s anonymity and even blaming an innocent client for potentially suspicious traffic. Here, the attacker’s idea is that during deanonymizations, the wrong encrypted identity is decrypted. Therefore, the relay could include an encrypted identity during two different relay extension handshakes. However, in this case, the hop key in the encrypted identity would not match the advertised hop key in the following encrypted identity . This mismatch can be undeniably blamed on the relay during deanonymizations.
Alternatively, the relay could advertise an encrypted identity , the hop key , relay identity , and its signatures on two different relay extensions. In this case, the next relay would, however, never receive a valid signature during the next step of the circuit extension. Hence, the circuit would never be established. In neither case is it possible for a malicious relay to undetectably reuse an encrypted identity to extend another client’s circuit.
8.1.2. Selectively Misbehaving Self-Hosted Relay
Similarly, malicious clients could attempt to cover up their traces by hosting their own relay. This relay would behave honestly for all other clients but incorrectly process the encrypted identity when a client creates a circuit through its own relay. In this case, the deanonymization process cannot identify the client. However, the deanonymization process ensures that the self-hosted relay is identified as misbehaving. Hence, the relay operator and consequently the client are identified as suspicious and will be further investigated by the LEA.
8.2. Deanonymization Attacks
While some adversaries may want to prevent deanonymizations through Seldom, others may aim to perform unauthorized deanonymizations of clients. In the following, we look at different attacker strategies, ranging from classical traffic analysis to collusion attacks, and how Seldom defends against these strategies.
8.2.1. Traffic Analysis Attacks
Traffic analysis is a common attack to deanonymize users of anonymity networks (Murdoch and Danezis, 2005; Panchenko et al., 2011). While Tor does not explicitly protect against global adversaries (Dingledine et al., 2004), i.e., an entity that intercepts all traffic, analyzing traffic between two or multiple nodes is often possible for adversaries. For our discussion, we focus particularly on the additional information that may be leaked about the client in Seldom compared to Tor. An observer intercepting the outgoing traffic of a client not using a bridge relay learns two things: the hop key of the circuit and the client’s intention to use Seldom. The hop key does not help in deanonymizing a circuit, and the intention to use the network is also leaked in Tor.
As the relays communicate through TLS channels, even during circuit establishment, no additional information is leaked besides the amount of transmitted data. However, this information is unlikely to give an adversary an advantage in detecting circuit extensions, as, usually, many circuits are multiplexed over a single link. In the end, such attacks can also be executed against traffic patterns of different websites (Panchenko et al., 2011, 2016; Hayes and Danezis, 2016), which is possible for both Tor and Seldom. All traffic leaving exit relays is either anonymized client traffic or encrypted deanonymization information. Neither leaks any additional information.
8.2.2. Mismatching Encrypted Identities
A malicious consortium member could potentially use partial deanonymizations to leak information about users. Therefore, the malicious member could set up a circuit that includes the threshold-encrypted identity of a targeted user. The consortium could be tricked to decrypt this identity, e.g., by marking the circuit as probe traffic.
However, such an attack would be detected after the consortium removes one layer of threshold encryption due to an invalid signature. The corresponding relay would be flagged as suspicious, but the malicious entity, as part of the consortium, would learn the identity of the preceding relay in the targeted circuit. The attacker could repeat this step and eventually learn the client’s identity. However, the consortium would notice the second iteration of this attack as the plaintexts are from different layers of the same circuit and thus clearly identify an attack from within the consortium. Thus, the client’s identity remains secure, as long as the attacker does not operate the entry relay of the targeted user. Additionally, the rest of the consortium and eventually also the public would learn about this one-off attack and could abandon the network or hold the consortium member accountable.
8.2.3. Collusion among Relays
@@
Seldomenables the exit relay to learn the encrypted client identity (cf. Figure 2) and indisputably link connections to it while preserving client anonymity (cf. Section 5). To show that this claim holds, we analyze which relays have access to which data and derive whether any relay has unintended access to client-identifying data. Here, we consider information that can be correlated within one circuit and across different circuits.
@@
During circuit establishment, the entry relay learns the client’s IP certificate, the hop keys and , as well as the signature created with . The certificate (and consequently the reuse of across multiple circuit establishments) contains no new information about the client, as the entry relay knows the client’s IP address and the current time.
@@
The hop key and, consequently, the signature created with , are unique per circuit and derived from a secure source of randomness. Thus, they do not allow any relay to gain information about clients from other circuits. The -th middle relay knows the hop keys and , the identity of the previous relay, and signatures by the previous relay and . All other information is threshold-encrypted by the -th relay. This nested encryption ensures that no earlier relay can match any information to deanonymize the client. In Tor, intermediary relays also know their direct neighbors. As is only shared with the preceding relay and only with the following relay, non-connected relays cannot notice that they are part of the same circuit. Crucially, the signature as the only link between these two keys is not shared with any other relay.
@@
8.3. Other Collusion Scenarios
Beyond the threats that Seldom explicitly defends against, additional collusion scenarios remain outside the scope of direct technical protection. Nonetheless, collusion between a CA and an entry relay as well as collusion among consortium members, is unlikely to materially compromise user anonymity.
8.3.1. CA and Entry Relay Collusion
A potential protocol-based attack arises from the possible collusion between a CA and an entry relay. Here, the malicious entities could create a fictional client with any IP address that is certified by the CA. If the circuit is then deanonymized for some reason, e.g., suspicious activities by the fictional client, the deanonymization process reveals the fake IP address. In this case, a powerful adversary could thus use Seldom to incriminate any person browsing the Internet. However, such incrimination would also be possible with other means by such adversaries. Most importantly, Seldom only removes anonymity and gives LEA hints for further investigations. IP addresses, allegedly used for illegal activities with Seldom or on other platforms, need to be supported by further evidence, as has been concluded by several courts (1; 35; 14). Thus, such collusion by powerful entities does not yield them capabilities they would not have without the existence of Seldom.
8.3.2. Collusion Within the Consortium
@@
To ensure that the risk of being covertly deanonymized in Seldom through collusion of a majority of consortium members and an exit relay is comparable to Tor, the consortium must be assembled accordingly. Here, it is important to recall that Tor does not fully protect the anonymity of its users but only decreases the chances that an attacker can learn a client’s identity (Arma, 2014). One attack to deanonymize a client in Tor is through the collusion of the three relays that make up a client’s circuit. Thus, Seldom’s deanonymizations should be more resilient to collusions than collusion of three randomly selected relays. As the exit relay is involved in the deanonymization attack in Tor and Seldom, we can remove it from the equation.
@@
Hence, we compare the risk of collusion in Seldom’s consortium to the risk of collusion of two randomly selected relays. There are multiple reasons why Seldom’s consortium is significantly less likely to collude than these two relays: while anyone can deploy massive numbers of relays, the identities of the consortium members are publicly known. Specifically, it would be necessary to convince a significant fraction of privacy advocates, e.g., some of the directory authorities, and foreign governments to partake in such an attack. Even attempting this could result in significant repercussions, which makes a successful collusion attack among the consortium members highly unlikely. We also designed Seldom in such a way that a higher number of members must participate in a decryption than have to accept a deanonymization vote, i.e., some disagreeing members have to accept the consensus and cooperate, to further minimize the risk of collusion.
@@ 9. Ethical and Technical Concerns
The safety-privacy dilemma remains particularly divisive, with LEA and privacy advocates maintaining seemingly fundamentally incompatible positions regarding a problem that requires some sort of compromise (Twenning and Baier, 2024). Fostering the understanding of technical means for enabling exceptional access without the dooming vision of privacy-invasive legislation or hidden backdoors for LEA, we address frequent concerns regarding Seldom’s design in the following.
@@
Q1 – Do we advertise for Seldom to replace Tor? No. Seldom relies on translucency and public oversight, which means that observed misbehavior by Seldom’s consortium must be met with revocations of these privileges, i.e., a migration of the user base back to Tor. In the far future, Seldom might attract a significantly larger user base than Tor and thus offer better anonymity even when accounting for selective deanonymization capabilities, especially if users start abandoning Tor. For Seldom’s security concept to work, there must, however, always exist a quick and easy way to migrate back to an anonymity network without deanonymization capabilities, even if temporarily deserted. This could be achieved with a browser that lets the users easily, and potentially collaboratively, toggle whether to use Tor or Seldom.
@@
Q2 – Can citizens trust the consortium? The underlying assumption for Seldom is that a majority of consortium members are trustworthy and value privacy by default, except for extraordinary situations that warrant deanonymization. Those members therefore have little incentive to partake in collusions, especially considering the high risk of such attempts being revealed to the public through the several translucency properties of Seldom. Further, Seldom protects against individual malicious consortium members. Namely, the majority of honest consortium members will faithfully delete their secret shares after encrypted identities reach the end of the mandated retention period or in extreme events, such as an imminent government takeover. This way, the consortium renders older encrypted identities unusable for malicious members even in the case of an anticipated dishonest majority gaining control over Seldom. However, maintaining transparency about the goals and non-goals of Seldom remains important to avoid creating false expectations about Seldom’s protection.
@@
Q3 – Can individual relays misbehave? Once a relay misbehaves and does not comply with deanonymization requests, they will become suspicious of the requesting LEA. Hence, they would reveal themselves as targets for further investigation. To reduce the number of misbehaving relays, Seldom can further use active probing (cf. Section 6.4) and remove any relays that misbehave.
@@
Q4 – Can investigations fail then? In exceptional cases, LEA will not be able to find the true root of harmful traffic using Seldom alone. For instance, misbehaving relays may cover up specific bad actors, or bad actors may establish circuits via compromised devices. However, all these cases lead LEA to a distinct location worth of further forensic investigation. Hence, Seldom will not always guarantee an immediate investigation success, but it provides crucial pointers to narrow down additional steps.
@@
Q5 – Can Seldom support hidden services? Technically, Seldom can integrate hidden services. Each introduction and rendezvous point would be associated with an encrypted identity, which could be deanonymized if deemed necessary. The consortium must, however, consider the increased privacy implications of agreeing to such a request. If a Seldom-like system were to be deployed, it should, however, be carefully considered whether hidden services should be supported or not.
@@
Q6 – Could Seldom interfere with Tor’s service? Initially, Seldom targets the population of privacy-conscious users who currently shy away from using Tor due to having negative associations by tying Tor to cyber-criminality. Only if Seldom is capable of establishing a reputation as the infrastructure of choice for privacy-seeking, law-abiding citizens, a potential user migration from Tor to Seldom could take place. As a positive effect, this would actively harm the achievable anonymity for cybercriminals, as they cannot hide among a mass of benign traffic. On the other hand, it becomes harder to abandon Seldom if a deserted Tor network offers little anonymity. To ensure that such a situation does not encourage extensive surveillance by the consortium, it must be easy for the user base to switch back to Tor at once, e.g., by offering a Tor and Seldom-compatible client that, by default, switches to Tor if Seldom is collectively deemed unsafe.
@@ 10. Conclusion
With this paper, we investigate the potential of selective deanonymizations for anonymity networks. We find that Seldom’s design barely impacts performance in comparison to Tor, while data overhead remains manageable. On the other hand, Seldom relies on a translucent trusted consortium and potentially encourages the migration of users from Tor, which would weaken the anonymity provided by Tor. With this paper we thus contribute to understanding the advantages and disadvantages of selective deanonymizations in anonymity networks like Seldom. These insights should encourage future research to mitigate drawbacks, and, most importantly, they should foster more open, objective, and technical discussions in inevitable future political debates about exceptional access.
@@ Appendix A Integration into Tor’s Handshake
We now discuss how Seldom integrates its oblivious authentication into Tor’s circuit establishment. Figure 11 depicts the extensions made to Tor’s circuit establishment. The data exchange of the original Tor handshake is compiled in the cell types, which are indicated in capital letters. Seldom extends the relevant cell types, i.e., Create2, Extend2, and Extended2 cells with the data necessary for the oblivious authentication. The clients and relays append the IP certificates or encrypted identities to Create2 cells. This data, however, does not always fit in a single cell. Therefore, the data is fragmented, and the first fragment is appended to the Create2 cell. All further fragments are sent in the newly added Create2_Addata cells. Seldom dispatches these cells immediately after transmitting Create2 cells. As soon as a relay receives the Create2 cell (not necessarily all Create2_Addata cells), it executes the Tor handshake and sends a Created2 cell back. After a relay has sent all Create2_Addata cells, it sends a standard Extended2 cell back to the client to inform it about the successful circuit extension. Once all Create2_Addata have been received by the extending relays, they perform all checks as explained in Section 5. Overall, Seldom’s oblivious authentication handshake can thus be integrated seamlessly into Tor’s circuit extension without introducing any additional round trips.
Appendix B Opaquely Retrieving IP Certificates